Problem 7-6 Bond Prices [L02] Sqeekers Co. issued 14-year bonds a year ago at a coupon rate of 8.6 percent. The bonds make semiannual payments and have a par value of $1,000. If the YTM on these bonds is 6.9 percent, what is the current bond price? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.) Current bond price
Answer #1

Number of periods = 13 * 2 = 26

Rate = 6.9% / 2 = 3.45%

Coupon = (0.086 * 1000) / 2 = 43

Current bond price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Current bond price = 43 * [1 - 1 / (1 + 0.0345)26] / 0.0345 + 1000 / (1 + 0.0345)26

Current bond price = 43 * [1 - 0.414006] / 0.0345 + 414.006472

Current bond price = 43 * 16.985333 + 414.006472

Current bond price = $1,144.38